OpenStudy (mathmath333):

k=1^2+2^2+3^2.......m^2. m<54. how many values of m exist such that k is divisible by 4?

OpenStudy (mathmath333):

a.10 b.11 c.12 d.none of these

OpenStudy (anonymous):

It should help to know that \[\large\sum_{i=1}^mi^2=\frac{m(m+1)(2m+1)}{6}=1+4+\cdots+m^2\]

OpenStudy (mathmath333):

what i have to do after finding k?

OpenStudy (anonymous):

Well \(k=\dfrac{m(m+1)(2m+1)}{6}\), and you want to see how many \(m\) allow \(k\) to be divisible by 4. My guess would be to expand, giving \[k=\frac{2m^3+3m^2+m}{6}=\frac{1}{3}m^3+\frac{1}{2}m^2+\frac{1}{6}m\] Then divide by 4: \[\frac{k}{4}=\frac{1}{12}m^3+\frac{1}{8}m^2+\frac{1}{24}m\] From here you have to find out which values of \(m\) make \(\dfrac{k}{3}\) a whole number. This would be any \(m\) whose cube is divisible by 12, square divisible by 8, and that is divisible by 24. One such value that works is \(m=24\), since \(\dfrac{m}{24}=1\), \(\dfrac{m^2}{8}=72\), and \(\dfrac{m^3}{12}=1152\).

OpenStudy (mathmath333):

u mean k/3 should be integer ?

OpenStudy (mathmath333):

48 is another value

OpenStudy (mathmath333):

@ganeshie8 are their other values than 24 and 48

ganeshie8 (ganeshie8):

Notice below : \(n^2\) leaves a remainder 0 when \(n\) is even (example 2^2 = 0 mod 4) \(n^2\) leaves a remainder 1 when \(n\) is odd (example 3^2 = 1 mod 4)

ganeshie8 (ganeshie8):

k=1^2 + 2^2 + 3^2 + ... +m^2 =1 + 0 + 1 + .... mod 4

ganeshie8 (ganeshie8):

So the sum is divisible by 4 when m is of form : 8t-1 or 8t

ganeshie8 (ganeshie8):

8t - 1 < 54 t < 55/8 (6 values for t) 8t < 54 t < 54/8 (6 values for t)

ganeshie8 (ganeshie8):

Overall there will be 12 sums divisible by 4

OpenStudy (mathmath333):

can u give link for comprehensive study of MOD

ganeshie8 (ganeshie8):

try these https://www.youtube.com/watch?v=SR3oLCYoh-I